are there any other results that of
 if f(x)f(1/x)=f(x)+f(1/x)
then,
f(x)= 1 \pm x^n
if any other such results please tell me

Grade:12th pass

1 Answers

Deepak Kumar Shringi
6 years ago
yeah there are results such that
f(x)+f(y)=f(xy) then f(x)=lnx
f(x)f(y)=f(x+y) then f(x)=ax